OpenStudy (anonymous):

Determine the slope and the y-intercept of the graph of the equation. -x + 10y = 40

OpenStudy (anonymous):

10y = x +40 y = 1/10x + 4 slope : 1/10 y intercept : 4

OpenStudy (anonymous):

- x + 10 y = 40 We will first write this equation in y = mx + b form , where m is slope and b is y-intercept Adding x to both the sides - x + x + 10 y = 40 + x 10 y = x + 40 Now dividing 10 to both the sides 10y/10 = x/10 + 40/10 y = (/10) x + 4 Now comparing it with y = mx + b , we get m = 1/10 and b = 4 So Slope (m) = 1/10 y-intercept (b) = 4 I hope this helps Thanks Abhishek Sinha